OpenStudy (anonymous):

Herman opened a credit card at a department store with an APR of 24.45% compounded semiannually. What is the APY on this credit card? PLEASE HELP!!!!

OpenStudy (amistre64):

\[(1+\frac{apr}{n})^n=1+apy\]

OpenStudy (amistre64):

in this case, semiannually means n=2

OpenStudy (anonymous):

can you explain it a little?

OpenStudy (amistre64):

APY is simple interest APR is compounded when does the APY equal the compunding of the APR?

OpenStudy (amistre64):

the setup i presented is the mathical equivalent of that statement ... solve for APY

OpenStudy (anonymous):

i dont know that is what im stuck on

OpenStudy (anonymous):

i dont get how to solve for APY?

OpenStudy (amistre64):

the setup is \[(1+\frac{apr}{n})^n=1+apy\] given that the apr is compounded semiannually (n=2), adn the apr = 24.45% \[(1+\frac{.2445}{2})^2=1+apy\] the rest is basic algebra .... subtract 1

OpenStudy (anonymous):

okay when you multiply the second 2 is it for everything in the parentesis?

OpenStudy (amistre64):

the exponent is applied to the paranthesis, yes

OpenStudy (anonymous):

 i dont get what to do after that?